# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v3

@@ Abstract
Accurate perception is essential for advancing autonomous driving and addressing safety challenges in modern transportation systems. Despite significant advancements in computer vision for object recognition, current perception methods still face difficulties in complex real-world traffic environments. Challenges such as physical occlusion and limited sensor field of view persist for individual vehicle systems. Cooperative Perception (CP) with Vehicle-to-Everything (V2X) technologies has emerged as a solution to overcome these obstacles and enhance driving automation systems. While some research has explored CP’s fundamental architecture and critical components, there remains a lack of comprehensive summaries of the latest innovations, particularly in the context of V2X communication technologies. To address this gap, this paper provides a comprehensive overview of the evolution of CP technologies, spanning from early explorations to recent developments, including advancements in V2X communication technologies. Additionally, a contemporary generic framework is also proposed to illustrate the V2X-based CP workflow, aiding in the structured understanding of CP system components. Furthermore, this paper categorizes prevailing V2X-based CP methodologies based on the critical issues they address. An extensive literature review is conducted within this taxonomy, evaluating existing datasets and simulators. Finally, open challenges and future directions in CP for autonomous driving are discussed by considering both perception and V2X communication advancements.
@@
Autonomous driving; connected and automated vehicles (CAVs); cooperative perception (CP); V2X Communication; edge computing; blockchain; 6G; digital twin; segmentation; object detection; multi-object tracking; data fusion; deep learning; artificial intelligence (AI).
@@ I Introduction
Perception plays a pivotal role in ensuring safety in autonomous driving, facilitating awareness of the surrounding road environment and other road participants over time, and utilizing various sensor modalities [1, 2, 3, 4, 5, 6, 7, 8]. Recent advancements in intelligent vehicle technology have led to significant progress in single-vehicle perception techniques across domains, such as 3D object identification using camera [9, 10, 11, 12, 13], LiDAR [14, 15, 16, 17, 18, 19, 20], radar [21, 22, 23, 24, 25, 26], and fusion [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], as well as 3D multi-object tracking employing LiDAR [42, 43, 44, 45, 46, 47], radar [48, 49, 50, 51], and fusion [52, 53, 54, 55, 56, 57, 58, 59], to name just a few. However, single-vehicle perception technologies still grapple with notable limitations. Sparse observations due to occlusion, restricted sensor Field of View (FoV), and sensor resolution remain primary challenges [60]. Additionally, susceptibility to sensor errors arising from adverse weather and hardware failures raises concerns about system robustness [61]. These bottlenecks constrain the potential and safety of automated vehicles in complex driving scenarios.
To address the above limitations, Cooperative Perception (CP) technology has emerged, focusing on interactions between multiple agents to overcome deficiencies in single-agent perception. The core technology of CP lies in the fusion of multi-agent sensor data, generating an accurate environmental representation, which is often seen in recent research on environmental monitoring [62], healthcare [63, 64, 65], and agriculture [66, 67, 68]. This approach has led to significant progress in the development of autonomous systems in the air [69, 70], maritime [71, 72], and on land [73, 74, 75]. CP in autonomous driving, facilitated by Vehicle-to-Everything (V2X) communications [76], offers a promising avenue to transcend the limitations of single-vehicle perception [77]. V2X communications enable Connected Autonomous Vehicles (CAVs) to enhance their perception capabilities through wireless information exchange, encompassing communication between vehicles (V2V) [78], roadside infrastructures (V2I) [79], and cloud networks (V2N) [80]. CP based on V2X significantly improves vehicle awareness on the road, addressing the challenges posed by high traffic volumes, pedestrian activity, and limited visibility. The grand vision is that V2X-based CP will become a fundamental element of future Intelligent Transportation Systems (ITSs).
(a) Without CP, each vehicle has a limited perception range. Thus, they cannot plan operations.
(b) With CP, each vehicle has a complete awareness of the traffic environment and is able to plan its operations.
Figure 1: Illustration of CP supported by V2X at an intersection.
(a) Without CP, the white hatchback has a blocked front view and thus potentially can have a fatal collision with the running boy.
(b) With CP, the white hatchback can see through the food truck, predicate the boy’s behaviour and avoid a potentially fatal collision.
@@
Paper Year Publication Survey Topic V2X Comm. The Taxonomy of Cooperative Perception System: Modules and Issues
EC BC DT 6G CAS PIS ACD APE CIF MH MD DRR LC DT PU TD DP S2R Sim DS
@@
[81] 2021 IEEE Network Feature level based cooperative object detection $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[82] 2022 IEEE IV V2I cooperative object detection and tracking $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$
@@
[83] 2022 IEEE T-ITS Cooperation issues in localization, map generation, object detection and tracking $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$
[84] 2023 IEEE ITS-M Multi-sensor fusion and cooperative perception $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$
@@
[85] 2023 IEEE ITS-M How to improve V2X cooperation efficiency, robustness, and safety $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\surd$
This survey 2023 -- Advancement in V2X communications supporting CP in CAVs. New taxonomy and modern generic framework of CP workflow. Challenges and future research opportunities in V2X-based CP. $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
@@
  • Note: V2X Comm.: V2X Communication, EC: Edge Computing, BC: Blockchain, DT: Digital Twin, CAS: Cooperative Agent Selection, PIS: Perception Information Selection, ACD: Alignment for Communication Delay, APE: Alignment for Pose Errors, CIF: Cooperative Information Fusion, MH: Model Heterogeneity, MD: Modality Distinctness, DRR: Data Rate Requirements, LC: Lossy Communication, DT: Data Trustworthiness, PU: Perception Uncertainties, TD: Task Discrepancy, DP: Data Privacy, S2R: Simulation to Reality, Sim: Simulator, DS: Dataset. The exact meaning of these items in this taxonomy will be explained systematically in the following sections.
CP in autonomous driving holds immense potential, promising an enhanced user experience, elevated road safety, improved air quality, diverse transportation applications, and a multitude of cutting-edge innovations [86]. Numerous additional factors must be considered to bring V2X CP into practical application, aligning with the foundational technologies underpinning V2X CP. For example, the associated data privacy [87, 88, 89] and ethical concerns [90], the associated development of smart cities [91], urban infrastructure [92, 93], the associated standardization, regulation, and policy from governments and regulatory bodies [94, 95, 96], the associated trust issue in human-system interaction [97], and the balance between excessive reliance and information inundation [98]. Thus, while the potential of V2X CP is profound, a concerted, multi-dimensional effort is requisite for its holistic integration into our daily lives.
An example of CP supported by V2X at an intersection is illustrated in Fig. 1. It is shown from Fig. 1a that each vehicle operates individually depending on its perception systems, and no collaborations exist. Thus, each vehicle has a limited understanding of the traffic environment and can only plan its operation in a short time frame. Note that the depicted perception range in this figure is purely illustrative. In reality, the sensing range varies depending on the types of sensors and weather conditions. However, due to obstructions and varying distances, autonomous vehicles face challenges such as unreliable perception of distant objects and complete blindness or noisy observation of objects blocked from their FoV [99]. In contrast, Fig. 1b demonstrates the scenario where vehicles collaborate to achieve CP. Therefore, CP enables CAVs to understand the traffic environment holistically, eliminating blind spots and ensuring complete awareness of the traffic situation for all vehicles involved. Fig. 2 provides a specific case extracted from Fig. 1, where a white hatchback, with its front view partially obstructed by a parked food truck. As shown in Fig. 2a, the white hatchback cannot directly observe a boy chasing a football near the road. This darting-out scenario presents a significant collision risk if the white hatchback fails to stop quickly. However, when CP is active among CAVs, as shown in Fig. 2b, the white hatchback can anticipate the boy’s movement, allowing itself to slow down in advance and prepare for an emergency stop, thus preventing a potentially fatal accident. CP proves instrumental in mitigating accidents and ensuring the safety of all parties in complex traffic environments.
Numerous scholarly investigations have been dedicated to CP for autonomous driving, substantiating its benefits. However, the emergence of CP introduces new challenges in driving automation, including data trustworthiness, malicious attacks, and limited communication bandwidth. With a growing number of methodologies focused on addressing these challenges, it is imperative to conduct a comprehensive survey of the advancements in this domain, analyze remaining gaps, and provide updated research problems and directions.
Several existing surveys have reviewed the progress in V2X cooperation technologies over the past few years. These papers have reviewed previous CP approaches for autonomous driving from various perspectives. For instance, a survey on feature-level-based cooperative object detection [81] outlined research outcomes in machine learning, feature map quantification, millimetre-wave communications, and vehicular edge computing. It discussed the selection of vital feature maps to mitigate data volume and communication latency. Another work [82] reviewed the research progress of infrastructure-based CP in object detection and tracking. However, these surveys often cover only specific aspects of CP, such as applications, feature-level cooperation, or V2I collaboration, hindering readers’ systematic understanding of existing cooperation technologies. An overview of algorithms and architectures for CP was provided in [83], emphasizing critical cooperation challenges related to localization, map generation, object detection, and tracking. Additionally, the survey presented in [84] elucidated multi-sensor fusion collaborative sensing techniques from the perspectives of V2V and V2I collaborative schemes, discussing challenges in CP for autonomous driving. Despite addressing prevalent CP concerns like communication delay, location errors, information fusion, and dataset collection, these surveys lacked an examination of CP research progress concerning practical implementation challenges. Consequently, a recent work [85] assessed collaborative modules and solutions based on cooperation efficiency, robustness, and safety considerations. Additionally, these surveys have yet to cover the importance of V2X communication technology. V2X communication technology is crucial for the success of CP in autonomous vehicles and intelligent transportation systems, as it serves as the backbone architecture for transferring information between CAVs in the network, ultimately leading to CP and achieving better decision-making for CAVs on the road. Table I compares our paper with these published surveys.
This work contributes to a comprehensive understanding of CP and its enabling V2X technologies and builds upon existing survey efforts. The key contributions are as follows:
  • Filling gaps in existing surveys by conducting a thorough review of the evolution of V2X CP from its inception in 2008 to state-of-the-art proposals until the end of 2023.
  • Discussing four key enabling V2X communication technologies: edge computing, blockchain, digital twins, and 6G communications.
Figure 3: Organization of this paper.
The structure of this paper is depicted in Fig. 3. Section II provides a brief historical overview of CP system development. Section III discusses recent advances in V2X communications, the foundational infrastructure for CP. Section IV presents a generic framework for comprehending the CP system systematically. Technical issues related to CP systems and associated datasets are summarized in Sections V and VI, respectively. Section VII explores open challenges and potential future directions, followed by the conclusion in Section VIII. A list of key acronyms and abbreviations used throughout the paper is given in Table II.

### v4

@@ Abstract
Achieving fully autonomous driving with heightened safety and efficiency depends on vehicle-to-everything (V2X) cooperative perception (CP), which allows vehicles to share perception data, thereby enhancing situational awareness and overcoming the limitations of the sensing ability of individual vehicles. V2X CP is crucial for extending perception range, improving accuracy, and strengthening the decision-making and control capabilities of autonomous vehicles in complex environments. This paper provides a comprehensive survey of recent advances in V2X CP, introducing mathematical models of CP processes across various collaboration strategies. We examine essential techniques for reliable perception sharing, including agent selection, data alignment, and fusion methods. Key issues are analyzed, such as agent and model heterogeneity, perception uncertainty, and the impact of V2X communication constraints like delays and data loss on CP effectiveness. To inspire further advancements in V2X CP, we outline promising avenues, including privacy-preserving artificial intelligence (AI), collaborative AI, and integrated sensing frameworks, as pathways to enhance CP capabilities.
@@
Autonomous driving; connected and automated vehicles (CAVs); cooperative perception (CP); V2X Communication; segmentation; object detection; multi-object tracking; data fusion; deep learning; artificial intelligence (AI).
@@ I Introduction
Perception plays a pivotal role in ensuring safety in autonomous driving, facilitating awareness of the surrounding road environment and other road participants over time, and utilizing various sensor modalities [1, 2, 3, 4, 5, 6, 7, 8]. Recent advancements in intelligent vehicle technology have led to significant progress in single-vehicle perception techniques across domains, such as 3D object detection using camera [9, 10, 11, 12, 13], LiDAR [14, 15, 16, 17, 18, 19], radar [20, 21, 22, 23, 24, 25], and fusion [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], as well as 3D multi-object tracking employing camera [42, 43, 44, 45], LiDAR [46, 47, 48, 49, 50, 51], radar [52, 53, 54, 55], and fusion [56, 57, 58, 59, 60, 61, 62, 63], to name just a few. However, single-vehicle perception technologies still grapple with notable limitations. Sparse observations due to occlusion, restricted sensor Field of View (FoV), and sensor resolution remain primary challenges [64]. Additionally, susceptibility to sensor errors arising from adverse weather and hardware failures raises concerns about system robustness [65]. These bottlenecks constrain the potential and safety of automated vehicles in complex driving scenarios.
(a) Without V2X CP, each vehicle has a limited perception range.
(b) With V2X CP, each vehicle gains an expanded awareness of its perception range.
Figure 1: An urban application scenario of V2X CP. The circles with different colours indicate the perception range, which is illustrative, as real sensing capabilities depend on factors such as sensor type and weather conditions.
(a) Without V2X CP, the white car has a blocked front view and thus potentially can have a fatal collision with the running boy.
(b) With V2X CP, the white car can detect the running boy due to the expanded perception range, enabling it to predict the boy’s behavior and avoid a potentially fatal collision.
@@
Cooperative Perception (CP) has been developed to mitigate the constraints associated with the perception capabilities of individual vehicles. Enabled by Vehicle-to-Everything (V2X) communication [66], CP allows Connected Autonomous Vehicles (CAVs) to exchange information with each other (V2V) [67], infrastructure (V2I) [68, 69], pedestrians (V2P) [70], and networks (V2N) [71]. This communication facilitates the sharing of crucial information, such as traffic conditions, road hazards, and signal status, extending beyond the perceptual range of individual vehicles [72]. Through this data exchange, CP enhances decision-making, enabling vehicles to better anticipate the movements of other road users, synchronize actions, and execute coordinated maneuvers, such as platooning and safe lane changes.
In full-stack autonomous driving, components like perception, localization, planning, control, and system integration collectively enable a vehicle to operate autonomously and safely. V2X CP significantly enhances this stack by sharing sensory and positional data via V2X communication between connected vehicles and infrastructure, thereby extending each vehicle’s awareness beyond the limits of its sensing range. Through V2X CP, the perception layer benefits from improved accuracy and safety, as vehicles can detect occluded or distant objects using shared data from other agents111“Agents” refer to autonomous entities within traffic environments, including vehicles and infrastructure equipped with sensors and communication systems, working together to enable cooperative perception. in the environment. In localization, V2X CP refines vehicle positioning by leveraging positional data from surrounding agents, which is especially valuable in areas with poor GPS signals. Planning also becomes more adaptive and safer, as V2X CP allows access to real-time data on nearby vehicles’ movements, enabling vehicles to make smoother, more informed maneuvers. Additionally, the enhanced situational awareness that V2X CP provides indirectly improves control precision and supports seamless system integration, reducing the vehicle’s reliance on any single sensor. Overall, V2X CP strengthens the robustness and adaptability of autonomous systems, laying a foundation for safer and more scalable deployment in real-world scenarios.
Fig. 1 illustrates an urban scenario of CP supported by V2X communication. In Fig. 1a, each vehicle operates independently, relying solely on its perception systems without collaboration, which restricts their understanding of the environment with a limited sensing range. Due to obstructions and varying distances, autonomous vehicles face challenges such as unreliable perception of distant objects and obstructed views [73]. In contrast, Fig. 1b shows vehicles and the infrastructure collaborating through wireless communication, leading to a holistic understanding of the environment and eliminating blind spots. A specific scenario in Fig. 2 demonstrates a white hatchback whose front view is obstructed by a parked food truck, preventing it from detecting a boy chasing a football near the road, as depicted in Fig. 2a. This presents a significant collision risk if the vehicle fails to stop promptly. However, with CP enabled, the blue CAV shares information about the boy’s movement with the white hatchback, allowing it to anticipate and decelerate preemptively, as shown in Fig. 2b, thereby avoiding a fatal accident.
While CP enables safer driving, numerous additional factors must be considered to bring V2X CP into practical application, aligning with the foundational technologies underpinning V2X CP. For example, the associated data privacy [74, 75, 76] and ethical concerns [77], the associated development of smart cities [78], urban infrastructure [79, 80], the associated standardization, regulation, and policy from governments and regulatory bodies [81, 82, 83], the associated trust issue in human-system interaction [84],
@@
Paper Year Publication Survey Topic The Taxonomy of Cooperative Perception System: Modules and Issues
CAS PIS PIC DSRC C-V2X LC CD APE CIF MH DH DS DP PU TD S2R Sim Data
@@
[85] 2021 IEEE Network Feature level based cooperative object detection. $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[86] 2022 IEEE IV V2I cooperative object detection and tracking. $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$
@@
[87] 2022 IEEE TITS Cooperation issues in localization, map generation, object detection and tracking. $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$
[88] 2023 IEEE MITS Multi-sensor fusion and CP. $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$
[89] 2024 IEEE TITS Fundamental components and core aspects of a CP system for enabling cooperative driving automation. $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$
@@
[90] 2023 IEEE MITS V2X cooperation efficiency, robustness, and safety. $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$
[91] 2023 -- The challenges of implementing CP within V2X systems and the practical applicability of different CP methodologies in V2X environments. $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\surd$
[92] 2024 IEEE TIV CP in intelligent transportation systems, with emphasizing road-to-vehicles collaboration at intersections. $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\surd$
[93] 2024 IEEE TITS The perception architecture and associated technologies within the Vehicle-Road-Cloud Integration System. $\times$ $\surd$ $\times$ $\surd$ $\surd$ $\times$ $\surd$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$
This survey 2024 -- The evolution of V2X CP technologies, key elements of V2X CP, and V2X CP solutions that tackle major issues in practical driving scenarios. $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
@@
  • Note: CAS: Cooperative Agent Selection, PIS: Perception Information Selection, PIC: Perception Information Compression, DSRC: Dedicated Short-range Communication, C-V2X: Cellular-based Vehicle-to-Everything, LC: Lossy Communication, CD: Communication Delays, APE: Alignment for Pose Errors, CIF: Cooperative Information Fusion, MH: Model Heterogeneity, DH: Data Heterogeneity, DS: Data Security, DP: Data Privacy, PU: Perception Uncertainties, TD: Task Discrepancy, S2R: Simulation to Reality, Sim: Simulator, Data: Dataset. The exact meaning of these items in this taxonomy will be explained systematically in the following sections.
and the balance between excessive reliance and information inundation [94]. Thus, while the potential of V2X CP is profound, a concerted, multi-dimensional effort is requisite for its holistic integration into our daily lives.
Although various approaches have been proposed to address these challenges, a comprehensive survey is needed to assess recent advancements, identify research gaps, and propose future directions. Some existing surveys often focus on specific aspects of CP, such as feature-level cooperative object detection [85] or infrastructure-based CP [86], but lack a systematic overview of V2X CP technologies. To fill this gap, the work in [87] provided an overview of CP algorithms, emphasizing challenges like localization and detection, while the review in [88] discussed communication delays and information fusion in multi-sensor V2V and V2I schemes. The survey in [89] explored CP system architectures and proposed a hierarchical framework for unifying CP scenarios. Despite covering common CP challenges, these surveys lacked an analysis of CP progress regarding practical implementation challenges.
Recent work [90] evaluated collaborative modules and solutions focusing on efficiency, robustness, and safety. Another survey [91] addressed practical challenges in implementing CP, such as latency, bandwidth constraints, and robustness, categorizing methodologies into early, intermediate, and late fusion. Environmental simulations were also conducted to assess mainstream CP approaches. A more recent survey [92] focused on CP at intersections, analyzing perception, communication, and downstream applications. However, these surveys overlook the critical role of V2X communication, which is essential for enabling effective CP in autonomous vehicles by facilitating information transfer and supporting decision-making. The work in [93] examined typical V2X communication technologies, including Dedicated Short-range Communication (DSRC) and Cellular-based V2X communication (C-V2X), and presented an overview of perception architectures and technologies within the Vehicle-Road-Cloud Integration System, encompassing single-node, multi-node, and vehicle-road-cloud cooperative perception frameworks. While it offers insights into these technologies, it does not explore specific algorithms and techniques in depth. This work contributes to a comprehensive understanding of CP and its enabling V2X technologies and builds upon existing survey efforts. Table I compares our paper with these surveys. The key contributions are as follows:
@@
DSRC Dedicated Short-range Communication
C-V2X Cellular-based Vehicle-to-Everything
IRS Intelligent Reflecting Surface
@@
  • Refer to caption
    Figure 3: Organization of this paper.
  • Filling gaps in existing surveys by conducting a thorough review of the evolution of V2X CP from its inception to state-of-the-art proposals.
  • Discussing established V2X communication technologies and the key impacts of real-world communication constraints on CP.
The structure of this paper is illustrated in Fig. 3. Section II presents the mathematical formulation of V2X CP. A brief historical overview of CP system development is provided in Section III. Section IV introduces a generic framework for systematically understanding the CP system. Practical issues related to CP systems and relevant datasets are discussed in Sections V and VI, respectively. Section VII compares the performance of various CP methods in offline testing and explores real-world implementations. Section VIII explores the challenges and potential future directions, and the paper concludes with Section IX. Table II lists key acronyms and abbreviations used throughout the paper.
II Mathematical Formulation of Cooperative Perception
This section begins with an overview of sensors and CP collaboration strategies, followed by a detailed mathematical formulation for early, intermediate, and late collaboration strategies.
II-A Overview
CP is built upon the shared data contributed by each participating agent. For each individual agent, raw observations are collected via onboard sensors, including cameras, LiDAR, radar, and, in some cases, ultrasonic sensors. Cameras provide rich visual information—such as color, texture, and shape—essential for object detection and scene interpretation [95]. LiDAR generates high-resolution 3D point clouds by measuring the return time of laser pulses, enabling precise distance and shape estimation critical for detailed mapping and obstacle detection [96]. Radar supplies range, speed, and movement data, effectively detecting objects at greater distances and under diverse weather conditions, making it especially valuable for tracking the velocity of surrounding vehicles. Although ultrasonic sensors are less commonly utilized in CP, they contribute valuable proximity data in close-range applications, such as parking and low-speed maneuvers.
CP strategies are categorized into three collaboration types: early, late, and intermediate collaboration, each defined by the level of data processing before sharing. Note that all types of sensor data can be utilized across any of these collaboration types. In early collaboration, raw sensor data is shared directly among agents, providing high-fidelity information but requiring substantial communication bandwidth and precise synchronization. This approach can be challenging in networks with limited bandwidth or high latency, leading to potential inconsistencies in data across agents. In late collaboration, only the final perception outputs, such as object classifications and location, are shared. This minimizes communication demands significantly but limits data granularity, potentially restricting the system’s ability to perform detailed analysis or adapt to complex, rapidly evolving scenarios. Intermediate collaboration strikes a balance by pre-processing sensor data to extract features before sharing them with other agents. This approach reduces data volume while retaining essential information, lowering bandwidth requirements. However, the effectiveness depends on each agent’s feature extraction accuracy, and variability in feature quality may lead to perception inconsistencies across agents.
A schematic overview of these collaboration schemes is provided in Fig. 4. For mathematical clarity, a scenario with $N$ collaborating agents is considered, where the local raw sensor data observed by the $i$-th agent at time step $t$ is denoted as $\boldsymbol{X}_{i}^{t}$. In the following subsections, the workflow and mathematical formulation of each collaboration scheme are elaborated in detail.
(a) Early
(b) Late
(c) Intermediate
Figure 4: Overview of collaboration schemes for autonomous driving: (a) Early collaboration; (b) Late collaboration; (c) Intermediate collaboration.
II-B Early Collaboration
As illustrated in Fig. 4a, early collaboration involves aggregating raw sensor observations collected by all participating agents, fostering a comprehensive perspective [97, 98, 99]. The $i$-th agent broadcasts its raw observation $\boldsymbol{X}_{i}^{t}$ to the other $N-1$ neighbouring agents and also receives the raw observation $\boldsymbol{X}_{j}^{t},j\in\{1,\cdots,N\},j\neq i$. After that, it performs data fusion on the received raw data with self-raw data and produces fused CP raw data $\boldsymbol{O}_{i}^{t}$. For simplicity, we assume that all agents use the same fusion, encoder, and decoder algorithms. Thus, the raw data fusion at the $i$-th agent can be expressed as:
$\boldsymbol{O}_{i}^{t}=\mathcal{F}_{\mathrm{EarlyFusion}}\left({\left\{\boldsymbol{X}_{k}^{t}\right\}}_{k=1}^{N}\right),$ (1)
where $\mathcal{F}_{\mathrm{EarlyFusion}}(\cdot)$ denotes the early fusion algorithm.
Then, the $i$-th agent uses an encoder $\mathcal{F}_{\mathrm{Encoder}}(\cdot)$ to extract high-level CP feature $\boldsymbol{\widetilde{F}}_{i}^{t}$ from its fused CP raw data $\boldsymbol{O}_{i}^{t}$:
$\boldsymbol{\widetilde{F}}_{i}^{t}=\mathcal{F}_{\mathrm{Encoder}}(\boldsymbol{O}_{i}^{t}).$ (2)
Finally, the $i$-th agent uses a decoder $\mathcal{F}_{\mathrm{Decoder}}(\cdot)$ to decode the CP feature $\boldsymbol{\widetilde{F}}_{i}^{t}$ and obtain the CP results $\widetilde{\boldsymbol{Y}}_{i}^{t}$:
$\widetilde{\boldsymbol{Y}}_{i}^{t}=\mathcal{F}_{\mathrm{Decoder}}\left(\boldsymbol{\widetilde{F}}_{i}^{t}\right).$ (3)
II-C Late Collaboration
In contrast to early collaboration, late collaboration involves exchanging individual perceptual outcomes, as depicted in Fig. 4b, requiring minimal data transmission [100].
In this case, the $i$-th agent independently processes its raw perception data $\boldsymbol{X}_{i}^{t}$, generate its feature data $\boldsymbol{F}_{i}^{t}$ and its final perception result $\boldsymbol{R}_{i}^{t}$, via (4) and (5) respectively:
$\boldsymbol{F}_{i}^{t}=\mathcal{F}_{\mathrm{Encoder}}\left(\boldsymbol{X}_{i}^{t}\right),$ (4)
$\boldsymbol{R}_{i}^{t}=\mathcal{F}_{\mathrm{Decoder}}\left(\boldsymbol{{F}}_{i}^{t}\right),$ (5)
where $\mathcal{F}_{\mathrm{Encoder}}(\cdot)$ and $\mathcal{F}_{\mathrm{Decoder}}(\cdot)$ represent the encoder and decoder algorithms, respectively.
After individual processing, the individual perception results from each agent are shared with all other agents. Then, for the $i$-th agent, it uses a late fusion algorithm $\mathcal{F}_{\mathrm{LateFusion}}(\cdot)$ to fuse the received individual perception result from all other $N-1$ agents and its own perception result to generate CP result $\boldsymbol{\widehat{Y}}_{i}^{t}$:
$\boldsymbol{\widehat{Y}}_{i}^{t}=\mathcal{F}_{\mathrm{LateFusion}}\left({\left\{\boldsymbol{R}_{k}^{t}\right\}}_{k=1}^{N}\right).$ (6)
II-D Intermediate Collaboration
In this case, as shown in Fig. 4c, each agent shared intermediate features rather than raw data or final perception results. Firstly, the $i$-th agent extracts its own feature data $\boldsymbol{F}_{i}^{t}$ from its raw perception data $\boldsymbol{X}_{i}^{t}$ as described in (4). Then, the extracted individual feature data are shared among all $N$ agents via V2X communication. After receiving shared feature data from other $N-1$ agents, the $i$-th agent use an intermediate fusion algorithm $\mathcal{F}_{\mathrm{InterFusion}}(\cdot)$ to generate the fused CP feature data $\boldsymbol{H}_{i}^{t}$:
$\boldsymbol{H}_{i}^{t}=\mathcal{F}_{\mathrm{InterFusion}}\left({\left\{\boldsymbol{F}_{k}^{t}\right\}}_{k=1}^{N}\right).$ (7)
Finally, the fused CP feature $\boldsymbol{H}_{i}^{t}$ is fed into the decoder $\mathcal{F}_{\mathrm{Decoder}}(\cdot)$ to obtain the final CP result ${\boldsymbol{Y}}_{i}^{t}$ at the $i$-th agent:
${\boldsymbol{Y}}_{i}^{t}=\mathcal{F}_{\mathrm{Decoder}}\left(\boldsymbol{H}_{i}^{t}\right).$ (8)